OpenStudy (anonymous):

write the polar equation in rectangular form:\[r={1\over{9cos\theta-3sin\theta}}\]

hartnn (hartnn):

do you know how to find polar form when r=acos +bsin ?

OpenStudy (anonymous):

\[x = rcos \Theta, y =r\sin \Theta \] \]

OpenStudy (anonymous):

yes @hartnn

hartnn (hartnn):

so, here, you can multiply and divide by 9cos + 3 sin

hartnn (hartnn):

so that denominator is of the form (a+b)(a-b)=..?

OpenStudy (anonymous):

that makes sense

hartnn (hartnn):

yeah, a constant in denominator and 9cos +3sin in numerator will brig the equation in the form of a cos +b sin

hartnn (hartnn):

hmm...actually you don't get a constant in the denominator...

hartnn (hartnn):

you have to multiply and divide by \(\sqrt{9^2+3^2}\) first.

OpenStudy (anonymous):

wait that just gives me\[r={{9cos\theta+3sin\theta}\over{9cos^2\theta-3sin^2\theta}}\]

OpenStudy (anonymous):

could i do this..

OpenStudy (anonymous):

@hartnn To me there seems to be an easier way...

hartnn (hartnn):

now if you multiply and divide by conjugate of denominator.... feel free to share... :)

OpenStudy (anonymous):

\[\begin{array}{rcl} r &=& \frac{1}{9\cos\theta - 3\sin\theta}\\ 9r\cos\theta - 3r\sin\theta &=& 1 \\ 9x-3y &=& 1 \end{array} \]

OpenStudy (anonymous):

\[r={1\over9cos\theta}-{1\over3sin\theta}\]\[r+{1\over3sin\theta}={1\over9cos\theta}\]\[9cos\theta\times r={3sin\theta}\]\[9r cos\theta\times r^2 = 3r sin \theta\]\[9x+x^2+y^2+=3y\]\[x^2+9x+{81\over4}+y^2-3y+{9\over4}={45\over2}\]\[(x+{9\over2})^2+(y-{3\over2})^2={45\over2}\]yes? :D

hartnn (hartnn):

:O so much simpler...how could i not think of this :P

OpenStudy (anonymous):

wow...i feel stupid....................

hartnn (hartnn):

i too :P

hartnn (hartnn):

good thought @wio

OpenStudy (anonymous):

that seems right and it is but im still unsure...idk why